graph: -2x+4y=8
i have 3 columns that are x, y and xy and requires it to be marked on a graph

Answer by math_tutor2020(3821)


This is one potential table that can be formed.
xy(x,y)
02(0,2)
23(2,3)
44(4,4)
65(6,5)
86(8,6)
107(10,7)


And this is the graph

Point A corresponds to (0,2)
Point B corresponds to (2,3)
and so on.

To get any row of that table, we pick a random x value to plug in.
Let's say we plugged in x = 0
Let's solve for y based on this x value.
-2x+4y = 8
-2(0)+4y = 8
0+4y = 8
4y = 8
y = 8/4
y = 2
So x = 0 leads to y = 2 to give the ordered pair (x,y) = (0,2)

I'll also show the steps for how x = 2 leads to y = 3
-2x+4y = 8
-2(2)+4y = 8
-4+4y = 8
4y = 8+4
4y = 12
y = 12/4
y = 3

I'll let you do the remaining steps. The table and graph above are there to check your work.

Technically you only need 2 points at minimum to plot the linear equation.
However I think it's good practice to plot more points. I'll leave it to you to decide how many points you want to do, or refer to your teacher for further instruction.

Side notes:
The x intercept is (-4,0) where the blue line crosses the x axis.
The y intercept is (0,2) where the blue line crosses the y axis.
The slope of the line is 1/2 meaning "go up 1, to the right 2".